Mechanical Vibration

2. problem 2 Find: (a) the euivalent mass with respect to x (b) the equivalent stiff ness with respect to a mi 000 m2
Answer #1

mig P — М, от, - bor bet ya aӨ Т - 40 6 - Kte - - ), а = — (1). б» (b+Т, a + + 0 = 0 bor mos me me tos mas M 2Fs m, DI - K2Substituts 69. (2) 2 (3) aco. 5.0 + (M268 +1620) 6+ (mia 8 + k, 90) a tk10=0 (dot man + miam 8 + (kast ki a 4K4) 2=0. Dea + k